You are planning to put a new digital flat TV on a wall that is 12 ft long and 9 ft
high. The digital TV has a diagonal of 72 inches. The length of the TV is twice
the width of the TV. How much of the wall will still need to be decorated
around the TV?

Answer by mananth(16946)

Area of wall = 9*12 = 108 ft^2
...
width of TV = x
Length = 2x
...
diagonal of TV = 72 "= 6 ft
..
x^2+(2x)^2=6^2
x^2+4x^2=36
5x^2=35
x^2=36/5
x = sqrt(36/5)
width = 2.68 ft.
length = 2.68*2=5.36 ft
Area of TV = 2.68*5.36
=14.36ft^2
..
Area of Wall - Area of TV= 108-14.36 = 93.63 ft^2
m.ananth@hotmail.ca